OpenStudy (anonymous):

solving a system of nonlinear equations: y=x^2 + 1 y=x+1

myininaya (myininaya):

y=x+1 and y=x^2+1 x+1=x^2+1 solve for x x^2+1-x-1=0 x^2-x=0 x(x-1)=0 x=0 or x=1 so y=x+1=0+1=1 y=x+1=1+1=2 so we have (0,1) and (1,2)

OpenStudy (anonymous):

thank you soo much! i was one step off.